OpenStudy (anonymous):

integral of e^(x^2)(2x^2+1)

OpenStudy (anonymous):

i tried substitution x^2=u ... and it seems that it could work but i'm not quite there yet

OpenStudy (dumbcow):

i don't know how to solve this with substitution but intuitively: let f(x) = e^(x^2) f'(x) = 2xe^(x^2) integral becomes f'(x)*x + f(x) this looks like product rule for derivatives d/dx x*f(x) = f'(x)*x + f(x) therefore anti-derivative is x*e^(x^2) +c

OpenStudy (anonymous):

WOW !!! very elegant solution !!! =]

OpenStudy (dumbcow):

thanks here is solution via wolfram http://www.wolframalpha.com/input/?i=integrate+%282x^2+%2B1%29e^%28x^2%29+dx notice it doesn't show any steps...